OpenStudy (reyrey8):

How many solutions does the equation have? 2x + 3 = 4x + 2

OpenStudy (reyrey8):

many solutions, no solutions, or 1 solution?

OpenStudy (thekingofcr):

one solution its 3

OpenStudy (reyrey8):

are you 100% sure

OpenStudy (thekingofcr):

pretty sure just solve it

OpenStudy (thekingofcr):

one sec let me make suurree

OpenStudy (reyrey8):

k

OpenStudy (thekingofcr):

2x + 3 = 4x + 2 3=2x+2 2x=1 yep almost positive

OpenStudy (reyrey8):

tanks <3

OpenStudy (thekingofcr):

was it right? i think it was

OpenStudy (reyrey8):

i hope